OpenStudy (cucacula):

solve the following equations for x and y : 5 x - y = 3 y² - 6x² = 25 what is the final answer?

OpenStudy (anonymous):

Are these two equations?

OpenStudy (cucacula):

yes only 2 equations . @zordoloom

OpenStudy (anonymous):

From the first equa. --> y = ... Plug it into the second one

OpenStudy (anonymous):

Okay, so one is a hyperbola, and the other is a straight line, Just set both euqal to 0 and simplify.

OpenStudy (anonymous):

you should get one x value to be 2 and one y value to equal 7. There is also two negative coordinate points.

OpenStudy (cucacula):

if i'm choose from equation one : 5 x - y = 3 y = 5x -3 for equation three ?? @Chlorophyll

OpenStudy (anonymous):

Yes, then plug it into the second one!

OpenStudy (anonymous):

x=2 x=-8/19 ,,,,,,, y=7 y=-97/19

OpenStudy (cucacula):

from equation one : 5 x - y = 3 y = 5x -3 ----> equation 3 sub 3 into 2 : y² - 6x² = 25 (3-5x)² - 6x² =25 (3-5x)(3-5x) -6x² =25 9-15x-15x+25x² -6x² =25 25x² -6x² -15x-15x+9-25 =0 19x² -30x-16 = 0 @zordoloom @Chlorophyll this correct??

OpenStudy (anonymous):

Yup, now solve it =)

OpenStudy (cucacula):

for x1 =2 and x2=-0.4211 is't true? @Chlorophyll

OpenStudy (anonymous):

Keep the result in fraction from!

OpenStudy (anonymous):

The plug x values back into the first equa. => y values

OpenStudy (cucacula):

sub 3 into x1 and x2 : y =5x-3 y=5(2) -3 y=10-3 y1=7 # y=5(x) -3 y=5(-0.4211)-3 y=-2.1055-3 y2=-5.1055 so x1= 2 x2=-0.4211 y1=7 y2=-5.1055 @Chlorophyll is't true answer?

OpenStudy (anonymous):

Match your answers to zordoloom's As I instructed you that, do NOT convert the x values into decimal numbers!

OpenStudy (cucacula):

how to convert x values not in decimal numbers? @Chlorophyll

OpenStudy (cucacula):

yeah i got already @Chlorophyll thanks you so much ,, :)